# Ensembling Portfolio Strategies for Long-Term Investments: A Distribution-Free Preference Framework for Decision-Making and Algorithms

### v1

@@ 4 Numerical experiments
The integrals of the combinatorial portfolios corresponding to (3.3) or (3.1) and (3.4) can be numerically approximated by Riemann sums that partition the simplices $\mathcal{B}^{N}$ and $\mathcal{B}^{k}$ into finitely discrete grids. Essentially, a finer partition can enhance the approximation, but it comes at the cost of increasing computational complexity, especially with high-dimensional simplices. Figure 2 provides an example from the experiment section, showcasing the numerical approximation of the investor’s combinatorial strategy using single-step size discretization for the 1-dimensional simplex. It illustrates how more discretization points and smaller step sizes can improve cumulative wealth. Additionally, we can also combine various step sizes and discretization points for further refinement. It is important to note that while the theoretically optimal constant combinations according to Definition 4 may not align with discretization points, fine partitions could ensure that numerical maximizers closely approximate theoretical values. However, it is worth mentioning that enhancing numerical approximation is not the focus of this experimental section.
Refer to captionRefer to caption
@@ 4.1 Data and testing strategies
We experimented with six diverse stocks over a 27-year span, from December 31, 1992, to December 31, 2019, covering $6798$ trading days. This period encapsulates major market events like the dot-com bubble, the 2008 financial crisis, and the COVID-19 pandemic, as illustrated in Figure 3. The selected stocks are blue-chip symbols, ranging across several sectors like industrials, energy, technology, and consumer defensive, traded on the NYSE and NASDAQ exchanges. They include Advanced Micro Devices, Inc. (AMD), The Boeing Company (BA), Honeywell International Inc. (HON), The Coca-Cola Company (KO), 3M Company (MMM), and Exxon Mobil Corporation (XOM). Their varied behaviors during historic events challenge prediction, making them ideal for testing our strategies. We used daily adjusted closing prices in USD provided by The Center for Research in Security Prices, LLC (CRSP).
@@
Market scenario. Let’s consider a market where many fund managers have different risk profiles for their investment strategies, and an investor can observe them as the components of a combinatorial strategy. We assume the component strategies are the sequences of Mean-Variance portfolios corresponding to the managers’ risk aversion levels $\beta\in\mathbb{R}_{+}$ as follows:
$b_{n}^{\beta}\coloneqq\operatorname*{argmax}_{b\in\mathcal{B}^{m}}\big{(}\beta\left\langle b,\hat{\mu}_{n}\right\rangle-\big{<}b^{\top}\hat{\varSigma}_{n},b\big{>}\big{)},\forall n,$
@@
$\hat{\mu}_{n}$ $\coloneqq\frac{{\displaystyle 1}}{J-1}\stackrel{{\scriptstyle[}}{{i}}=n-J]{n-1}{\sum}x_{i}\text{ \,\,and\,\, }\hat{\varSigma}_{n}\left(i,j\right)\coloneqq{\displaystyle\frac{1}{J-1}}\stackrel{{\scriptstyle[}}{{h}}=n-J]{n-1}{\sum}\big{(}x_{h,i}-\hat{\mu}_{n,i}\big{)}\big{(}x_{h,j}-\hat{\mu}_{n,j}\big{)},\text{ $\forall i,j$$\in\left[6\right],\forall n>J,$}$
@@
    The aforementioned Mean-Variance strategies (M-V), denoted as $\big{(}b_{n}^{\beta}\big{)}$ with various coefficients $\beta\in\mathbb{R}_{+}$, which serve as the component strategies for all the tested multi-strategies.
@@
    We propose an accelerated variant for the combinatorial strategy by using a time-varying density instead of the time-invariant uniform one over the simplex $\mathcal{B}^{k}$. Namely, we introduce the Universal combination on the Winners (UC-W) as an UC strategy, where the uniform density $\mu_{n}$ has support on $B_{n}\subset\mathcal{B}^{k}$ at period $n\geq 2$, defined as:
@@
    Accordingly, the Universal combination on the Losers (UC-L) is defined as:
@@
The experiments illustrate the evolution of cumulative wealths of the testing strategies through several graphics. Specifically, we investigate whether they exhibit a strict preference over the baseline strategies by eventually exceeding all the M-V component strategies. Additionally, we report additional performance measures for the testing strategies, including the final cumulative wealth, average growth rate, average return, and the empirical Sharpe ratio.
@@ 4.2 Experiments with small scales of component strategies
$\lim_{n\to+\infty}\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{\text{FL}}\big{)}\big{)}=\lim_{n\to+\infty}\frac{1}{n}\left(\log S_{N+1}\big{(}\ddot{b}_{N+1}\big{)}-\log S_{N+1}\big{(}b_{N+1}^{\text{FL}}\big{)}\right)=0.\vspace{0.25ex}$
However, the mechanisms of the FL and WAE strategies inherently constrain their growth rates, being bounded by the baseline strategy; consequently, despite the possibility of an indifferent preference with the latter, the cumulative wealth evolution of the two former strategies poorly exhibits. These experiments emphasize that attaining the same asymptotic growth rate as the best component strategy does not guarantee a multi-strategy not being the worst performer.
Table 1 presents performance metrics for the strategies involved, while Figure 5 displays the final cumulative wealth against the numerical constant combination. The Sharpe ratios of the multi-strategies do not surpass the best among the component M-V strategies. However, the UC strategies offer the most benefit as they achieve the highest final cumulative wealths in the first three experiments, with relatively small tradeoffs in terms of the Sharpe ratios. In contrast, all M-V strategies that yield higher final wealth and mean of return experience significant decreases in Sharpe ratios. These favorable performances of the UC strategies are attained through approximations with respect to the numerical combinations $\lambda$. Although their performances could potentially improve using a finer grid, Figure 5 demonstrates that such marginal improvement may not be worthwhile, as the computation becomes infeasible for a large scale of component strategies.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
The first row of Figure 6 illustrates the growth rate differences between the benchmark strategy $\big{(}b_{n}^{*}\big{)}$ and the UC strategy $\big{(}b_{n}^{\text{UC}}\big{)}$, as well as between the benchmark strategy $\big{(}b_{n}^{*}\big{)}$, which is also numerically computed, and the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$ of all relating M-V strategies. The Combination criterion is satisfied in the first three experiments but not in the remainder, as $\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{*}\big{)}\big{)}\to 0$ as $n$ increases. In Experiment 1, the M-V strategy $\big{(}b_{n}^{0.005}\big{)}$ almost always exceeds the M-V strategy $\big{(}b_{n}^{1}\big{)}$, but the UC strategy is able to exceed the former. Conversely, in Experiment 4, the UC strategy never exceeds the best M-V strategy $\big{(}b_{n}^{0.002}\big{)}$, which almost always exceeds the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$. In the two remaining experiments, the UC strategy is capable of exceeding the best M-V strategies, where one component strategy experiences strong fluctuations while the other two experience stable growth over time in Experiment 2, and five comparable component strategies continuously compete for the leading positions over time in Experiment 4.
@@
The second row of Figure 6 depicts the upper bound for the difference $\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}^{\text{UC}}\big{)}$ as established in Proposition 4, between the two numerically computed strategies. Furthermore, in order to fully comprehend the behavior of the benchmark strategy $\big{(}b_{n}^{*}\big{)}$, the first row of Figure 7 exhibits the best numerically constant combinations over time with respect to different discretization step sizes in each experiment. The first three graphics indicate the presence of only two non-zero components of the time-variant maximizer $\lambda_{n}$, while the last graphic shows only one non-zero component. This is due to the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$ consistently dominating the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$ in the sense of Proposition 3, which explains the dissatisfaction of the Combination criterion in Experiment 4. In contrast, the second row of Figure 7 illustrates the proportions in the UC strategy that the investor allocates to each component M-V strategy over time; every M-V strategy is consistently invested throughout.
@@
Universally combinatorial strategy with acceleration. The results of the first four experiments emphasize the necessity of accelerating the UC strategy to enhance the converging speed of its growth rate to that of the benchmark strategy or even surpassing it. Since the UC strategy allocates capital to all component strategies over time, it also invests in the component strategies with inferior performances. Particularly, the underperformance of the UC strategy compared to the significantly better FL strategy in Experiment 4 justifies our motivation to incorporate the mechanism of the latter into the former. To address this issue, the proposed UC-L and UC-W strategies aim to reduce the amount of capital invested in underperforming component strategies. However, these accelerated strategies do not have a theoretical guarantee like the original UC strategy. Therefore, we conduct four additional experiments to empirically examine their performances. The UC strategy will be tested alongside UC-L and UC-W in Experiments 5, 6, 7, and 8. In accordance with the definitions of UC-L and UC-W strategies, we consider the time-variant sets $B_{n}$ and $\bar{B}_{n}$ of constant combinations, which are defined in (4.2) and (4.3), in the top and bottom $30\%$ and $50\%$ of cumulative wealth from the previous time period $n-1$ among all combinations in the simplex $\mathcal{B}^{k}$. Only in Experiment 8, we further consider an additional UC-W strategy with the top $5\%$ of combinations. The performances of these strategies are presented in Figure 8 and Table 2.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
The results presented in Figure 8 align with our expectations. Since the FL strategy works in Experiment 4, it results in the UC-W ($5\%$) strategy demonstrating the most favorable evolution of cumulative wealth compared to the other tested strategies in the same scenario in Experiment 8. This can be attributed to the consistent exceeding of the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$ over the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$, which makes the strategies corresponding to a constant combination closer to the best one $\lambda_{n}=(1,0)$ tend to yield a higher cumulative wealth. Consequently, the UC-W strategy performs better with a small subset $B_{n}$ including the point $(1,0)$, while the UC-L strategy, with a small subset $\bar{B}_{n}$ including the point $(0,1)$, leads to a decline in cumulative wealth. However, in the other three experiments, the UC-W strategies underperform, whereas the UC-L strategies show significant improvement. A possible explanation for this behavior could be similar to the argument explaining the poor performance of the FL strategy in the corresponding scenarios. When there are more strongly fluctuating component strategies among the involved M-V strategies, relying solely on constant combination strategies that previously led in terms of cumulative wealth becomes more insecure. Additionally, Table 2 reaffirms a characteristic inherent in the UC-W and UC-L strategies derived from the UC strategy, namely, the strategy with a higher final cumulative wealth tends to have a small decrease in Sharpe ratio.
@@
The acceleration results in a noticeable change in the capital allocation, as depicted in the first row of 9. In Experiment 8, the UC-W ($5\%$) strategy primarily allocates capital to the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$, which closely aligns with the optimal constant combination $(1,0)$ over time. Conversely, in the remaining three experiments, the UC-L ($30\%$) strategy exhibits significant fluctuations in the allocated proportions among the involved component M-V strategies over time, in contrast to the gradual changes observed in the proportions of the UC strategy. These substantial fluctuations in allocated proportions subsequently impact the differences in growth rate between the benchmark strategy and the accelerated strategies, as shown in the second row of Figure 9. The first three graphs indicate that the growth rate $W_{n}\big{(}b_{n}^{\text{UC-L}}\big{)}$ is not bounded by $W_{n}\big{(}b_{n}^{*}\big{)}$, whereas $W_{n}\big{(}b_{n}^{\text{UC-W}}\big{)}$ converges to the upper bound $W_{n}\big{(}b_{n}^{*}\big{)}$ in the final graph. The accelerated UC-L strategy in Experiments 6 and 7 quickly surpasses the benchmark strategy, which could be caused by the number of component M-V strategies that exhibit significant fluctuations in terms of cumulative wealth evolution. In Experiment 5, the consistent exceeding of the M-V strategy $\big{(}b_{n}^{0.005}\big{)}$ over the M-V strategy $\big{(}b_{n}^{1}\big{)}$ enables the UC-W strategy, which allocates its capital to a small set of constant combination strategies that previously led in cumulative wealth, to perform better than the UC-L strategies most of the time. However, as the M-V strategy $\big{(}b_{n}^{1}\big{)}$ experiences stronger fluctuations after about $5000$ time periods, the UC-L strategies gain momentum and eventually surpass the UC-W strategies. In Experiments 6 and 7, characterized by significant fluctuations in at least one M-V strategy and frequent changes in the leading position, the UC-L strategy, which consistently allocates its capital to a small set of constant combination strategies that do not lead in cumulative wealth but may provide stability in return, quickly surpasses the benchmark strategy.

### v2

@@ 4 Numerical experiments
The integrals of the combinatorial portfolios corresponding to (3.3) or (3.1) and (3.4) can be numerically approximated by Riemann sums that partition the simplices $\mathcal{B}^{N}$ and $\mathcal{B}^{k}$ into finitely discrete grids. Essentially, a finer partition can enhance the approximation, but it comes at the cost of increasing computational complexity, especially with high-dimensional simplices. Figure 2 provides an example from the experiment section, showcasing the numerical approximation of the investor’s combinatorial strategy using single-step size discretization for the 1-dimensional simplex. It illustrates how more discretization points and smaller step sizes can improve cumulative wealth. Additionally, various step sizes and discretization points can be combined for further refinement. It is important to note that while the theoretically optimal constant combinations according to Definition 4 may not align with discretization points, fine partitions can ensure that numerical maximizers closely approximate theoretical values. However, enhancing numerical approximation is not the primary focus of this experimental section.
@@ 4.1 Data and testing strategies
We experiment with six diverse stocks over a 27-year span, from December 31, 1992, to December 31, 2019, covering $6798$ trading days. This period encapsulates major market events like the dot-com bubble, the 2008 financial crisis, and pre-COVID-19 pandemic, as illustrated in Figure 3. The selected stocks are blue-chip symbols, ranging across several sectors like industrials, energy, technology, and consumer defensive, traded on the NYSE and NASDAQ exchanges. They include Advanced Micro Devices, Inc. (AMD), The Boeing Company (BA), Honeywell International Inc. (HON), The Coca-Cola Company (KO), 3M Company (MMM), and Exxon Mobil Corporation (XOM). Their varied behaviors during historic events challenge prediction, making them ideal for testing our strategies. We used daily adjusted closing prices in USD provided by The Center for Research in Security Prices, LLC (CRSP).
@@
Market scenario. Let’s consider a market where many fund managers have different risk profiles for their investment strategies, and an investor can observe them as the components of a combinatorial strategy. We consider the sequences of Mean-Variance portfolios corresponding to the managers’ risk aversion levels $\alpha\in\mathbb{R}_{+}$, which are defined as follows, as component strategies:
$b_{n}^{\alpha}\coloneqq\operatorname*{argmax}_{b\in\mathcal{B}^{m}}\big{(}\alpha\left\langle b,\hat{\mu}_{n}\right\rangle-\big{<}b,\hat{\varSigma}_{n}b\big{>}\big{)},\forall n,$
@@
$\hat{\mu}_{n}$ $\coloneqq\frac{{\displaystyle 1}}{J}\stackrel{{\scriptstyle[}}{{i}}=n-J]{n-1}{\sum}x_{i}\text{ \,\,and\,\, }\hat{\varSigma}_{n}\left(i,j\right)\coloneqq{\displaystyle\frac{1}{J-1}}\stackrel{{\scriptstyle[}}{{h}}=n-J]{n-1}{\sum}\big{(}x_{h,i}-\hat{\mu}_{n,i}\big{)}\big{(}x_{h,j}-\hat{\mu}_{n,j}\big{)},\text{ $\forall i,j$$\in\left[6\right],\forall n>J,$}$
@@
    The aforementioned Mean-Variance strategies (M-V), denoted as $\big{(}b_{n}^{\alpha}\big{)}$ with various coefficients $\alpha\in\mathbb{R}_{+}$, which serve as the component strategies for all the tested multi-strategies.
@@
    The accelerated variant of the combinatorial strategy uses a time-varying density in place of the time-invariant uniform one over the simplex $\mathcal{B}^{k}$. Namely, we introduce the Universal combination on the Winners (UC-W) as a UC strategy, where the uniform density $\mu_{n}$ has support on $B_{n}\subset\mathcal{B}^{k}$ at period $n\geq 2$, defined as follows:
@@
    Accordingly, the Universal combination on the Losers (UC-L) is formed using the subsets:
@@
The experiments illustrate the evolution of cumulative wealths of the testing strategies through several graphics. Specifically, we investigate whether they exhibit a strict preference over the baseline strategies by eventually exceeding all the M-V component strategies. Additionally, we report common performance measures for the testing strategies, including their final cumulative wealth, growth rate, average return (i.e., empirical expected return), and empirical Sharpe ratio.
@@ 4.2 Experiments with small scales of component strategies
$\lim_{n\to\infty}\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{\text{FL}}\big{)}\big{)}=\lim_{n\to\infty}\frac{1}{n}\big{(}\log S_{N+1}\big{(}\ddot{b}_{N+1}\big{)}-\log S_{N+1}\big{(}b_{N+1}^{\text{FL}}\big{)}\big{)}=0.\vspace{0.25ex}$
However, the mechanisms of the FL and WAE strategies inherently constrain their growth rates, being bounded by the baseline strategy. Consequently, despite the possibility of an indifferent preference with the latter, the cumulative wealth evolutions of the two former strategies are worse. These experiments emphasize that attaining the same asymptotic growth rate as the best component strategy does not guarantee a multi-strategy not being the worst performer.
Table 1 presents performance metrics for the strategies involved, while Figure 5 displays the final cumulative wealth against the numerical constant combination. The Sharpe ratios of the multi-strategies do not surpass the best among the component M-V strategies. However, the UC strategies offer the most benefit as they achieve the highest final cumulative wealths in the first three experiments, with relatively small tradeoffs in terms of the Sharpe ratios. In contrast, all M-V strategies that yield higher final wealth and mean of return experience significant decreases in Sharpe ratios. These favorable performances of the UC strategies are attained through approximations with respect to the numerical combinations $\lambda$. Although their performances could potentially improve with a finer grid, Figure 5 demonstrates that such marginal improvement is not worthwhile, as the computation may become infeasible for a large scale of component strategies.
@@
The first row of Figure 6 illustrates the growth rate differences between the numerically computed benchmark strategy $\big{(}b_{n}^{*}\big{)}$ and the UC strategy $\big{(}b_{n}^{\text{UC}}\big{)}$, as well as between the benchmark strategy and the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$ of all related M-V strategies. The Combination criterion is satisfied in the first three experiments but not in the remainder, as $\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{*}\big{)}\big{)}\to 0$ as $n$ increases. In Experiment 1, the M-V strategy $\big{(}b_{n}^{0.005}\big{)}$ almost always exceeds the M-V strategy $\big{(}b_{n}^{1}\big{)}$, but the UC strategy is able to exceed the former. Conversely, in Experiment 4, the UC strategy never exceeds the best M-V strategy $\big{(}b_{n}^{0.002}\big{)}$, which almost always exceeds the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$. In the two remaining experiments, the UC strategy is capable of exceeding the best M-V strategies, where one component strategy experiences strong fluctuations while the other two experience stable growth over time in Experiment 2, and five comparable component strategies continuously compete for the leading positions over time in Experiment 4.
@@
The second row of Figure 6 depicts the upper bound for the difference between the two numerically computed strategies, $\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}^{\text{UC}}\big{)}$, as established in Proposition 4. Furthermore, in order to fully comprehend the behavior of the benchmark strategy $\big{(}b_{n}^{*}\big{)}$, the first row of Figure 7 exhibits the best numerically constant combinations over time with respect to different discretization step sizes in each experiment. The first three graphics indicate the presence of only two non-zero components of the time-variant maximizer $\lambda_{n}$, while the last graphic shows only one non-zero component. This is due to the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$ consistently dominating the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$ in the sense of Proposition 3, which explains the violation of the Combination criterion in Experiment 4. In contrast, the second row of Figure 7 illustrates the proportions in the UC strategy that the investor allocates to each component M-V strategy over time, showing that all of them are consistently invested in throughout.
@@
Universally combinatorial strategy with acceleration. The results of the first four experiments emphasize the necessity of accelerating the UC strategy to enhance the converging speed of its growth rate to that of the benchmark strategy or even surpassing it. Since the UC strategy allocates capital to all component strategies over time, it includes unnecessarily those with inferior performance. Particularly, the underperformance of the UC strategy compared to the significantly better FL strategy in Experiment 4 justifies our motivation to incorporate the mechanism of the latter into the former. To address this issue, the proposed UC-L and UC-W strategies aim to reduce the amount of capital invested in underperforming component strategies. However, unlike the original UC strategy, these accelerated strategies do not come with a theoretical guarantee. Therefore, the UC strategy will be tested alongside both of its variants UC-L and UC-W in Experiments 5, 6, 7, and 8 to empirically examine their performances. In accordance with the definitions of UC-L and UC-W strategies, we consider the time-variant sets $B_{n}$ and $\bar{B}_{n}$ of constant combinations, which are defined in (4.2) and (4.3), in the top and bottom $30\%$ and $50\%$ of cumulative wealth from the previous time period $n-1$ among all combinations in the simplex $\mathcal{B}^{k}$. Only in Experiment 8, we further consider an additional UC-W strategy with the top $5\%$ of combinations. The performances of these strategies are presented in Figure 8 and Table 2.
@@
The results presented in Figure 8 align with our expectations. Since the FL strategy works in Experiment 4, it results in the UC-W ($5\%$) strategy demonstrating the most favorable evolution of cumulative wealth compared to the other tested strategies in the same scenario in Experiment 8. This can be attributed to the almost always exceedance of the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$ over the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$, which causes a combinatorial strategy corresponding to a constant combination closer to the best one, $\lambda_{n}=(1,0)$ for all $n$, to yield higher cumulative wealth. Consequently, the UC-W strategy performs better with a small subset $B_{n}$ that includes the point $(1,0)$, while the UC-L strategy with a small subset $\bar{B}_{n}$ that includes the point $(0,1)$ leads to a decline in cumulative wealth. However, in the other three experiments, the UC-W strategies underperform, whereas the UC-L strategies show significant improvement. A possible explanation for this behavior is similar to the argument explaining the poor performance of the FL strategy in the corresponding scenarios. When strongly fluctuating strategies are present among the component M-V strategies, relying solely on constant combination strategies that previously led in terms of cumulative wealth becomes less reliable. Additionally, Table 2 reaffirms a characteristic inherent in the UC-W and UC-L strategies derived from the UC strategy, namely, that a strategy with higher final cumulative wealth tends to produce a slight decrease in its Sharpe ratio.
@@
The acceleration results in a noticeable change in the capital allocation, as depicted in the first row of 9. In Experiment 8, the UC-W ($5\%$) strategy primarily allocates capital to the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$, which closely aligns with the optimal constant combination $(1,0)$ over time. Conversely, in the remaining three experiments, the UC-L ($30\%$) strategy exhibits significant fluctuations in the allocated proportions among the involved component M-V strategies over time, in contrast to the gradual changes observed in the proportions of the UC strategy. These substantial fluctuations in allocated proportions subsequently impact the differences in growth rate between the benchmark strategy and the accelerated strategies, as shown in the second row of Figure 9. The first three graphs indicate that the growth rate $W_{n}\big{(}b_{n}^{\text{UC-L}}\big{)}$ is not bounded by $W_{n}\big{(}b_{n}^{*}\big{)}$, whereas $W_{n}\big{(}b_{n}^{\text{UC-W}}\big{)}$ approaches the upper bound $W_{n}\big{(}b_{n}^{*}\big{)}$ in the final graph. The accelerated UC-L strategy in Experiments 6 and 7 quickly surpasses the benchmark strategy, likely due to the number of component M-V strategies that exhibit significant fluctuations in cumulative wealth evolution. In Experiment 5, the consistent exceeding of the M-V strategy $\big{(}b_{n}^{0.005}\big{)}$ over the M-V strategy $\big{(}b_{n}^{1}\big{)}$ enables the UC-W strategies, which allocate their capital to small sets of constant combination strategies that previously led in cumulative wealth, to perform better than the UC-L strategies most of the time. However, as the M-V strategy $\big{(}b_{n}^{1}\big{)}$ experiences stronger fluctuations after about $5000$ time periods, the UC-L strategies gain momentum and eventually surpass the UC-W strategies. In Experiments 6 and 7, characterized by significant fluctuations in at least one M-V strategy and frequent changes in the leading position, the UC-L strategy, which consistently allocates its capital to a small set of constant combination strategies that do not lead in cumulative wealth but provide stability in returns, quickly surpasses the benchmark strategy.